# Stable Mean Teacher for Semi-supervised Video Action Detection

### v1

@@ Abstract
In this work, we focus on semi-supervised learning for video action detection. Video action detection requires spatio-temporal localization in addition to classification and limited amount of labels make the model prone to unreliable predictions. We present Stable Mean Teacher, a simple end-to-end student-teacher based framework which benefits from improved and temporally consistent pseudo labels. It relies on a novel ErrOr Recovery (EoR) module which learns from students’ mistakes on labeled samples and transfer this to the teacher to improve psuedo-labels for unlabeled samples. Moreover, existing spatio-temporal losses does not take temporal coherency into account and are prone to temporal inconsistencies. To overcome this, we present Difference of Pixels (DoP), a simple and novel constraint focused on temporal consistency which leads to coherent temporal detections. We evaluate our approach on four different spatio-temporal detection benchmarks, UCF101-24, JHMDB21, AVA and Youtube-VOS. Our approach outperforms the supervised baselines for action detection by an average margin of 23.5% on UCF101-24, 16% on JHMDB21, and, 3.3% on AVA. Using merely 10% and 20% of data, it provides a competitive performance compared to the supervised baseline trained on 100% annotations on UCF101-24 and JHMDB21 respectively. We further evaluate its effectiveness on AVA for scaling to large-scale datasets and Youtube-VOS for video object segmentation demonstrating its generalization capability to other tasks in the video domain. We will make the code and models publicly available.
@@ Introduction
Semi-supervised learning (SSL) is an active research area with two prominent approaches: iterative proxy-label (Rizve et al. 2020) and consistency based methods (Tarvainen and Valpola 2017). Iterative proxy-label methods, although effective, are not suitable for the video domain due to their lengthy training cycles. On the other hand, consistency-based approaches offer end-to-end solutions, requiring only a single pass through the dataset for training. While most of the existing research in this area has focused on image classification (Rasmus et al. 2015; Tarvainen and Valpola 2017; Sajjadi, Javanmardi, and Tasdizen 2016; Laine and Aila 2017) and object detection (Xu et al. 2021; Chen et al. 2022; Liu, Ma, and Kira 2022), limited efforts have been made in the video domain with works only focusing on classification (Jing et al. 2021; Xu et al. 2022). We also observe that Mean Teacher (Tarvainen and Valpola 2017) based approaches have demonstrated superior performance among consistency-based methods. Building upon the success of student-teacher learning in image domain, we extend it to video domain for spatio-temporal detection tasks.
@@
Video action detection Video action detection comprises two tasks: action classification and spatio-temporal localization. Some of the initial attempts to solve this problem are based image-based object detectors such as RCNN (Ren et al. 2015) and DETR (Carion et al. 2020), where detection at frame-level is used for video-level activity classification (Yang et al. 2019; Hou, Chen, and Shah 2017; Yang, Gao, and Nevatia 2017; Zhao et al. 2022; Chen et al. 2023; Ntinou, Sanchez, and Tzimiropoulos 2024). Most approaches involve two-stages, where localization is performed using a region proposal network which is classified into activities in the second stage (Gkioxari et al. 2018; Yang et al. 2019; Hou, Chen, and Shah 2017; Yang, Gao, and Nevatia 2017). Recently, some encoder-decoder based approaches have been developed on CNN (Duarte, Rawat, and Shah 2018) and transformer-based (Zhao et al. 2022; Wu et al. 2023; Chen et al. 2023; Ntinou, Sanchez, and Tzimiropoulos 2024) backbones which simplify the two-stage video action detection process. However, transformer-based backbones are complex and heavy, involving multiple modules. In a recent work (Kumar and Rawat 2022), the authors further simplify VideoCapsuleNet (Duarte, Rawat, and Shah 2018) to reduce computation cost with minor performance trade-off. In this work, we make use of this optimized approach as our base model for video action detection.
@@
Different from all these, we focus on videos where the temporal dimension adds more complexity to the problem. There are some recent works focusing on videos, but they are limited to video classification (Jing et al. 2021; Singh et al. 2021; Xiao et al. 2022; Xu et al. 2022) and temporal action localization (Ji, Cao, and Niebles 2019; Wang et al. 2021; Nag et al. 2022) where per frame dense spatio-temporal annotations is not required. We focus on video action detection, which requires spatio-temporal localization on every frame of the video in addition to video level class predictions. More recently, a PI-based consistency approach (Kumar and Rawat 2022) has been explored for semi-supervised video action detection. Different from this, we propose a Mean Teacher based approach adapted for video activity detection task which achieves better performance.

### v2

@@ Abstract
In this work, we focus on semi-supervised learning for video action detection. Video action detection requires spatio-temporal localization in addition to classification and limited amount of labels make the model prone to unreliable predictions. We present Stable Mean Teacher, a simple end-to-end student-teacher based framework which benefits from improved and temporally consistent pseudo labels. It relies on a novel ErrOr Recovery (EoR) module which learns from students’ mistakes on labeled samples and transfer this to the teacher to improve psuedo-labels for unlabeled samples. Moreover, existing spatio-temporal losses does not take temporal coherency into account and are prone to temporal inconsistencies. To overcome this, we present Difference of Pixels (DoP), a simple and novel constraint focused on temporal consistency which leads to coherent temporal detections. We evaluate our approach on four different spatio-temporal detection benchmarks, UCF101-24, JHMDB21, AVA and Youtube-VOS. Our approach outperforms the supervised baselines for action detection by an average margin of 23.5% on UCF101-24, 16% on JHMDB21, and, 3.3% on AVA. Using merely 10% and 20% of data, it provides a competitive performance compared to the supervised baseline trained on 100% annotations on UCF101-24 and JHMDB21 respectively. We further evaluate its effectiveness on AVA for scaling to large-scale datasets and Youtube-VOS for video object segmentation demonstrating its generalization capability to other tasks in the video domain. Code and models are publicly available at:
Codehttps://github.com/AKASH2907/stable-mean-teacher
Modelshttps://huggingface.co/akashkumar29/stable-mean-teacher
@@ Introduction
Semi-supervised learning (SSL) is an active research area with two prominent approaches: iterative proxy-label (Rizve et al. 2020) and consistency based methods (Tarvainen and Valpola 2017). Iterative proxy-label methods, although effective, are not suitable for the video domain due to their lengthy training cycles. On the other hand, consistency-based approaches offer end-to-end solutions, requiring only a single pass through the dataset for training. While most of the existing research in this area has focused on image classification (Rasmus et al. 2015; Tarvainen and Valpola 2017; Sajjadi, Javanmardi, and Tasdizen 2016; Laine and Aila 2017) and object detection (Xu et al. 2021; Chen et al. 2022; Liu, Ma, and Kira 2022), limited efforts have been made in the video domain with works only focusing on classification (Jing et al. 2021; Xu et al. 2022; Kumar et al. 2023). We also observe that Mean Teacher (Tarvainen and Valpola 2017) based approaches have demonstrated superior performance among consistency-based methods. Building upon the success of student-teacher learning in image domain, we extend it to video domain for spatio-temporal detection tasks.
@@
Video action detection Video action detection comprises two tasks: action classification and spatio-temporal localization. Some of the initial attempts to solve this problem are based image-based object detectors such as RCNN (Ren et al. 2015) and DETR (Carion et al. 2020), where detection at frame-level is used for video-level activity classification (Yang et al. 2019; Hou, Chen, and Shah 2017; Yang, Gao, and Nevatia 2017; Dave et al. 2022; Zhao et al. 2022; Chen et al. 2023; Ntinou, Sanchez, and Tzimiropoulos 2024). Most approaches involve two-stages, where localization is performed using a region proposal network which is classified into activities in the second stage (Gkioxari et al. 2018; Yang et al. 2019; Hou, Chen, and Shah 2017; Yang, Gao, and Nevatia 2017). Recently, some encoder-decoder based approaches have been developed on CNN (Duarte, Rawat, and Shah 2018) and transformer-based (Zhao et al. 2022; Wu et al. 2023; Chen et al. 2023; Ntinou, Sanchez, and Tzimiropoulos 2024) backbones which simplify the two-stage video action detection process. However, transformer-based backbones are complex and heavy, involving multiple modules. In a recent work (Kumar and Rawat 2022), the authors further simplify VideoCapsuleNet (Duarte, Rawat, and Shah 2018) to reduce computation cost with minor performance trade-off. In this work, we make use of this optimized approach as our base model for video action detection.
@@
Different from all these, we focus on videos where the temporal dimension adds more complexity to the problem. There are some recent works focusing on videos, but they are limited to video classification (Jing et al. 2021; Singh et al. 2021; Xiao et al. 2022; Xu et al. 2022) and temporal action localization (Ji, Cao, and Niebles 2019; Wang et al. 2021; Nag et al. 2022) where per frame dense spatio-temporal annotations is not required. We focus on video action detection, which requires spatio-temporal localization on every frame of the video in addition to video level class predictions. More recently, a PI-based consistency approach (Kumar and Rawat 2022; Singh et al. 2024) has been explored for semi-supervised video action detection. Different from this, we propose a Mean Teacher based approach adapted for video activity detection task which achieves better performance.
@@ Conclusion
Stable Mean Teacher for Semi-supervised Video Action Detection
Here, we go through some additional quantitative and qualitative results, extra ablation studies and, implementation details. Section I provides quantitative results on baseline setup. Section II provides comparison at different percentages against previous semi-supervised approaches. Section III provides some more discussions on JHMDB21 and UCF101-24. Section IV discuses some implementation details mentioned in main paper. Section V shows extra qualitative results.
Baseline Setup
We setup the baseline spatio-temporal mean teacher (STMT) and show it’s comparison to baseline Mean teacher and our proposed approach. We observe an improvement of approx. 5% f-mAP@0.5 on UCF101-24 and 4.5% on JHMDB-21. At v-mAP@0.5, the gain 4-5% on both the datasets. The final proposed approach further boost the score on top of baseline STMT model by a margin of 1.6% on UCF101-24 and 7% on JHMDB-21 at f-mAP@0.5. At v-mAP@0.5, the gain is 1.8% on UCF101-24 and 8% on JHMDB-21.
UCF101-24 JHMDB21
Method Annot. f-mAP v-mAP Annot. f-mAP v-mAP
% 0.5 0.2 0.5 % 0.5 0.2 0.5
Baseline Mean Teacher(Tarvainen and Valpola 2017) 10% 67.3 92.7 70.5 20% 56.3 88.8 52.8
ST-MeanTeacher (Ours) 10% 72.2 95.1 74.5 20% 61.8 94.7 62.0
Stable Mean Teacher (Ours) 10% 73.9 95.8 76.3 20% 69.8 98.8 70.7
Supervised baseline 10% 53.5 77.2 49.7 20% 55.7 93.9 52.4
Table 5: Analysis on Baseline Spatio-temporal mean teacher (STMT). STMT vs Stable Mean Teacher.
Comparison with previous Semi-supervised approaches
We extended the Table 1 in main paper to have detailed comparison on more semi-supervised approaches at multiple labeled percentages on UCF101-24 and JHMDB21 in Tables 6 and 7 respectively.
We show that our proposed approach outperforms previous semi-supervised approaches on multiple thresholds.
5% 8% 10% 15%
Method Annot. f-mAP v-mAP Annot. f-mAP v-mAP
f@0.5 v@0.5 f@0.5 v@0.5 f@0.5 v@0.5 f@0.5 v@0.5
Le et al. (Lee et al. 2013) 52.5 - 56.4 - 59.3 58.3 64.9 -
Jeong et al. (Jeong et al. 2021) 53.2 51.5 57.8 61.2 60.2 64.0 63.9 65.2
Kumar et al. (Kumar and Rawat 2022) 59.1 58.8 64.1 64.6 65.2 66.7 68.3 70.2
Tarvainen et al. (Tarvainen and Valpola 2017) 58.4 59.2 66.2 69.3 67.3 70.5 71.5 74.7
Ours 66.2 68.2 72.1 73.8 73.9 76.3 75.5 77.9
Supervised baseline 37.5 30.1 42.6 39.4 53.5 49.7 58.8 55.6
Table 6: Comparison with previous state-of-the art approaches on semi-supervised learning on UCF101-24 on multiple labeled subset.
10% 15% 20% 25%
f@0.5 v@0.5 f@0.5 v@0.5 f@0.5 v@0.5 f@0.5 v@0.5
Jeong et al. (Jeong et al. 2021) 48.5 46.2 54.3 51.8 57.8 57.0 59.5 58.1
Kumar et al. (Kumar and Rawat 2022) 45.8 41.8 52.1 50.1 59.1 58.7 61.4 60.5
Tarvainen et al. (Tarvainen and Valpola 2017) 47.6 44.3 60.9 60.4 61.8 62.0 65.4 66.0
Ours 54.2 50.3 66.8 66.5 69.8 70.7 71.3 70.9
Supervised baseline 43.7 37.7 50.2 48.6 55.7 52.4 60.4 59.3
Table 7: Comparison with previous state-of-the art approaches on semi-supervised learning on JHMDB21 on multiple labeled subset.
Extra Discussions
Firstly, we extend the same discussions as mentioned in main paper to UCF101-24 dataset on static vs dynamic, network architecture and performance improvement in low-label regime.
Static vs Dynamic scenes
Activities can be categorized into two sub-classes based on the background namely static, where background is constant, and dynamic, when the background changes. To give an overview, example of few classes categorized as static and dynamic, in UCF101-24, are basketball, golf swing, rope climbing, soccer juggling, tennis swing and basketball dunk, biking, cliff diving, diving, skateboarding respectively. Dynamic is a challenging situation since the actor and background both is changing in each frame. It’s also evident from lower results as compared to Static actions from Table 8(a) on supervised settings. Our proposed solution makes a 18.8% improvement on static and 23.3% on dynamic actions at f-mAP@0.5. This shows that proposed approach is able to localize actor in challenging situations as well.
Analysis on ErrOr Recovery (EoR) Module
Here, we analyze the effect of different architectures of EoR network. We replace the EoR network with a 2D version and saw that the performance degrade by a small margin (Table 8(b)) which supports 3D based version generates better spatio-temporal pseudo labels similar to results shown on JHMDB21.
Effectiveness of approach in low-label regime
Similar to JHMDB21, we look into performance gain at multiple ratios of labeled and unlabeled on UCF101-24. Looking into Table 8(c), we see that we have huge performance gain much more evident than JHMDB21 at lower percentages. The gain at 5% and 8% is almost double the gain at 10%.
Burn-in vs end-to-end?
Some recent approaches have shown the benefit of pre-training on labeled set for model initialization (Liu et al. 2021). In this experiment, we analyzed the effect of burn-in on proposed Stable Mean Teacher and pre-trained the model on labeled set and use it for combined training on labeled and unlabeled sets. We observe that there was no substantial gain to Stable Mean Teacher with burn-in weights.
Comparison on more thresholds
Here, we extend ablation table on EoR Module and compare the improvement by sub modules on JHMDB21 dataset. We show results on three more thresholds at 0.3, 0.4 and 0.6. Table 9 shows that proposed sub-modules helps in improvement more for higher thresholds.
(a) Static vs Dynamic.
Category Static Dynamic
20% sup. 67.5 36.9
20% semi 86.3 60.2
100% sup. 87.4 65.1
(b) Network Architecture.
Arch. f@0.5 v@0.5
Base 73.4 75.8
2D 73.6 76.0
3D 73.9 76.3
(c) Low label regime
Sup. Semi. $\uparrow$ %age
5% 37.5 66.2 76.5%
8% 42.6 72.1 69.2%
10% 53.5 73.9 38.1%
Table 8: Analysis on UCF101-24 on multiple factors. f@0.5 and v@0.5 denotes f-mAP@0.5 and v-mAP@0.5 respectively. Tables 8(a) and 8(c) shows performance on f-mAP@0.5.
L2 EoR DoP f-mAP v-mAP
@0.3 @0.4 @0.5 @0.6 @0.3 @0.4 @0.5 @0.6
83.6 74.2 61.8 46.4 92.8 79.7 62.0 38.2
89.2 81.2 68.3 49.6 96.8 85.2 68.1 42.8
84.4 75.0 62.9 46.5 90.7 82.2 64.5 39.3
91.9 84.6 69.8 50.5 95.9 85.6 70.7 42.9
Table 9: Ablation study on sub-modules at multiple thresholds. L2: Base mean teacher model with L2 as spatio-temporal localization loss, EoR: EoR network, and DoP: Difference of pixel constraint.
Figure 7: This figure shows the top 5 classes which has the most improvement on v-mAP@0.5 on our proposed semi-supervised approach compared to the supervised counterpart on JHMDB21 dataset.
Figure 8: This figure shows the top 5 classes which has the most improvement on v-mAP@0.5 on our proposed semi-supervised approach compared to the supervised counterpart on UCF101-24 dataset.
Strong Augmentations
Type Probability Random Value Explanation
Contrast 0.7 0.8 Random uniform selection between [0.6, 1.4)
Hue 0.7 0.05 Random uniform selection between [-0.1, 0.1)
Brightness 0.7 0.9 Random uniform selection between [0.6, 1.4)
Saturation 0.7 0.7 Random uniform selection between [0.6, 1.4)
Grayscale 0.6 - -
Gaussian Blur 0.5 $\sigma_{x}$=0.1, $\sigma_{y}$=2.0 Kernel size=(3, 3)
Weak + Strong Augmentation
Horizontal Flip 0.5 - -
Table 10: Details about selection of random parameters for spatial augmentations.
Figure 9: Visualization of augmentations: This figure shows the original clip and augmented clip from UCF101 and JHMDB21 dataset respectively.
Classwise Performance Analysis
In this study, we deep diver into the f-mAP and v-mAP of different classes. Here, we discuss performance at specific threshold of 0.5. From the figures 7, we show that the classes with the most improvement with our Stable Mean Teacher approach. The classes with improvement on f-mAP@0.5 and v-mAP@0.5 are brush_hair, kick_ball, sit, walk, wave and brush_hair, jump, sit, throw, walk respectively. Some of these classes have very fast motion. Most improvement on those classes shows that our approach is more robust to motion changes and the predictions are more temporally coherent.
We extend this analysis to UCF101-24 dataset as well. From Fig. 8, classes with the most gain are CliffDiving, Diving, HorseRiding, Skijet, Surfing, and, Diving, Skijet, Skateboarding, Surfing, DogWalking. A major boost in Diving and Surfing also corroborates our claim that Stable Mean Teacher is less susceptible large motion changes and also small objects.
Figure 10: Qualitative results - Case - I - Boundary Refinement In this scenario, we can see the Ours could even separate out the instance of two legs separately which shows that the precise error from EoR model helps in refinement for fine-grained details. The predictions are even better than the 100% supervised model.
Figure 11: Qualitative results - Case - II - Noise Suppression In this scenario, we can see the Ours is able to suppress the background noise more better. The detaching of EoR module from main model helps this procedure. Otherwise, the mispredictions gets enhanced.
Figure 12: Qualitative results - Case - III - Noise Suppression + Boundary Refinement In this scenario, model is able to do both getting rid of noise and refining the boundary at the same time. Even 100% supervised model fails at it.
Figure 13: Qualitative results - Case - IV - Temporal Mask Coherency In this scenario, we show that Ours not only helps to localize the actor spatially but the temporal coherency of mask is also maintained in case of large displacement/motion.
Implementation Details
We go through architecture, data augmentation and training details in depth here.
EoR Architecture
In our work, we use a modified version of UNet 3D architecture. It is a simple extension of it’s 2D version where 2D Convolution block is replaced by a 3D convolution block and the upsample mode is trilinear instead of bilinear. Original UNet 3D model has a lot of trainable parameters and to reduce the extra overhead of trainable parameters, we reduce the depth in our EoR Module. Original model has 5 channels depth, and the variation in depth goes like this, $32\rightarrow 64\rightarrow 128\rightarrow 256\rightarrow 512\rightarrow 256\rightarrow 128\rightarrow 64\rightarrow 32$. In our case, we reduce the number of channels. EoR architecture have this variation, $16\rightarrow 32\rightarrow 64\rightarrow 128\rightarrow 64\rightarrow 32\rightarrow 16$. This brings down the number of trainable parameters to approximately 1.1M. We also changed the EoR model with various depth and compared the performance.
Data Augmentation Details
We study both spatial and temporal augmentations to generate weak and strong views. First, video is passed through temporal augmenter block, which temporally augments the video frames. After temporal augmentation, the video passes through the spatial augmenter block. Augmenting in this sequence makes the process computationally efficient as for spatial augmentation we only perform augmentation of required frames instead of augmenting all the video frames. The strong augmentation includes random crop, gaussian blur, horizontal flip, grayscale, hue, saturation, brightness, and contrast, whereas, the weak augmentation includes random crop and horizontal flip.
We break into two: 1) Spatial: For a weak augmented view, only random horizontal flip is applied. On the other hand, for a strong view, all types of augmentations are applied. To measure the consistency between student and teacher prediction, if the frames of weak augmented video is flipped, then, correspondingly frames of strong augmented video is flipped, i.e., geometrical transformation is maintained. 2) Temporal: This augmentation is also similar for both teacher and student. This is because we need to calculate localization consistency across each frame. We determine this hyperparameter empirically. One out of three temporal augmentations is chosen randomly equal probability.
Training Details
Some extra training details: The weight is gradually ramped up till 15 epochs and kept consistent after that. We use Adam optimizer with an initial learning rate set to 0.0001. Next, we discuss the augmenter setup to generate the two augmented views. In our work, we use different set of augmentations for weak and strong augmented view. For weak, we use Random Horizontal Flip, and, for strong, we use Color Jitter, Grayscale and Gaussian Blur. In Table 10, we show the random probability parameters with which these are augmentations are applied on the video. We have not search for best hyperparameter settings for data augmentation in our work. In Fig. 9, an example of spatial augmentation only is shown for two videos, one from UCF101 and one from JHMDB21.
Qualitative Analysis
Figures 10 - 13 show couple of more examples showing analyzing the model output qualitatively between different settings. In all figures, GT means ground truth. Successive rows shows predicted localization maps for 100% and 20% fully supervised trained model and Ours.